For the beam given below, determine the resultant internal loading acting on the cross section at C. Use the following assumption: The beam is weightless. Support A is a roller, and support B is a pin. The cross section C is 1 m away from A. (Tips: Internal loading includes normal force, shear force, and moment). . 100 kN 50 kN/m ol B 2m 1m 2m
Answer #1

25 kN/m 12.5 kN/m Mc Nc Vc 0.5 m 1 m

According to an above figure, we have

w / (1 m) = (50 kN/m) / (2 m)

w = (25 kN/m2) (1 m)

w = 25 kN/m

Magnitude of the resultant of distributed load is equal to the area under the loading curve (triangle) and acts through the centroid of this area.

Then, we have

F = (1/2) (25 kN/m) (1 m)

F = 12.5 kN

Which acts [(1 m) / 2] = 0.5 m from C.

For an equilibrium conditions, we have

sum Fx = 0 ;   - NC = 0

NC = 0

sum Fy = 0 ; VC - (12.5 kN) = 0

VC = 12.5 kN

Therefore, the resultant internal loading acting on the cross section at C which will be given as -

sum MC = 0 ;   - MC - (12.5 kN) (0.5 m) = 0

- MC - (6.25 kN.m) = 0

MC = - 6.25 kN.m

{ The negative sign indicates that MC acts in the opposite direction